OpenStudy (anonymous):

How many Solutions Does The System Have? x=-4y-2 3x+12y=-6 Please Help, I Don't Understand How To Do Any of This Thank You =)

OpenStudy (anonymous):

3(-4y-2)+12y=-6 -12y-6+12y=6 0=0 none verify..?

OpenStudy (anonymous):

Cool! thanks The Multiple Choice Answer D.) No Solution Is Correct. Thank You Lol